A partnership has the following capital balances:

Henry (60% of gains and losses) $ 52,000
Thomas (20%) 60,000
Catherine (20%) 124,000

Anne is going to invest $80,000 into the business to acquire a 30 percent ownership interest. Goodwill is to be recorded. What will be Anne’s beginning capital balance?

Multiple Choice

  • $70,800

  • $94,800

  • $80,000

  • $101,143
Answer #1

answer: $101,143

Calculation:

80,000 + goodwill = 0.3(236,000 + 80,000 + goodwill)


80,000 + goodwill = 94,800 + 0.3(goodwill)

goodwill - 0.3(goodwill) = 14,800

0.7(goodwill) = 14,800

goodwill = 21,143

capital balance = 21,143 + 80,000 = $101,143